An IQ test is designed so that the mean is 100 and the standard deviation is 17 for the population of normal adults. Find the sample size necessary to estimate the mean IQ score of statistics students such that it can be said with 95​% confidence that the sample mean is within 7 IQ points of the true mean. Assume that standard deviation is 17 and determine the required sample size using technology. Then determine if this is a reasonable sample size for a real world calculation.

The required sample size is____

​(Round up to the nearest​ integer.)
Answer #1

Solution :

Given that,

standard deviation = = 17

margin of error = E = 7

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.960

Sample size = n = ((Z/2 * ) / E)2

= ((1.960 * 17 ) / 7 )2

= 22.6576

= 23

Sample size = 23